One crew can seal a parking lot in 8 hours and another in 10 hours. How long will it take to seal the parking lot if the two crews work together?
Answer by ptaylor(2198)


Let x=amount of time needed when both crews are working together
So both crews working together works at the rate of 1/x parking lot per hr
First crew works at the rate of 1/8 parking lot per hr
The other crew works at the rate of 1/10 parking lot per hr
So our equation to solve is:
1/8 + 1/10=1/x multiply each term by 40x
5x+4x=40
9x=40
x=4.4444444 hrs-------------time needed when both crews are working together
CK
(1/8)*4.444 +(1/10)*4.4444=1 (1 parking lot, that is)
0.555555555 +0.4444444444=1
0.99999999999~~~~~1 close enough

Hope this helps-----ptaylor